# DINO in the Room: Leveraging 2D Foundation Models for 3D Segmentation

### v1

@@ 3.2 Distillation
While injecting DINOv2 features directly into the 3D backbone can significantly boost segmentation performance, relying on calibrated images at inference time may be restrictive in certain real-world scenarios. To address situations where only 3D data can be used at test time, we propose a distillation scheme (D-DITR) to transfer knowledge from the frozen 2D encoder into a pure 3D model.
@@ 4.1 Datasets
We evaluate our method on well-established 3D indoor and outdoor segmentation benchmarks. For indoor segmentation, we select ScanNet [8], ScanNet200 [44], and S3DIS [1]. All three datasets consist of colored point clouds and their corresponding RGB-D frames. For outdoor segmentation, we select nuScenes [4], SemanticKITTI [3], and Waymo [50]. nuScenes provides six cameras with combined $360^{\circ}$ surround coverage and captures point clouds using a 32-beam LiDAR sensor. Waymo offers five cameras covering the front and sides, and uses a 64-beam main LiDAR along with four short-range LiDAR sensors, resulting in a dense point cloud. SemanticKITTI also uses a 64-beam LiDAR, resulting in a density similar to Waymo, but includes only a single front-facing camera.
@@ 4.2 Implementation Details
We use PTv3 [58] as our 3D backbone, as it is the state-of-the-art model for 3D semantic segmentation across both indoor and outdoor datasets. Unless stated otherwise, we retain the default architecture and learning hyperparameters of PTv3 to ensure fair comparison; additional details are provided in the supplementary material. For all datasets, we adhere to standard evaluation protocols and metrics. Reproducing PTv3 results on the S3DIS [^1] https://github.com/Pointcept/Pointcept/issues/154 and SemanticKITTI [^2] https://github.com/Pointcept/Pointcept/issues/186 datasets has been notoriously hard for the community. Therefore, we also report our reproduced results as a reference.
@@
During both training and inference on outdoor scenes, we use all available camera views: one for SemanticKITTI, five for Waymo, and six for nuScenes. These camera views appear in a consistent orientation with respect to the LiDAR sensor and can be assumed always to be present for all scans. For the indoor datasets, the number of RGB-D frames per scene can vary tremendously and always surpasses the number of frames we can effectively process with the available GPU memory. Therefore, unless specified otherwise, we select 10 uniformly sampled, temporally equidistant frames from each RGB-D video during inference, while randomly sampling 10 views per scene during training to enhance data diversity. Since DINOv2 is pretrained at a maximum resolution of $518\times 518$, we resize input images to maintain a similar amount of patches while preserving aspect ratio [54]. Further details are provided in the supplementary material.
4.3 Quantitative Results
Method ScanNet ScanNet200 S3DIS
Val Test Val Test Area5
@@ 4.3 Quantitative Results
DITR 80.5 79.7 41.2 44.9 74.1
Table 1: Indoor semantic segmentation results (mIoU). We obtain state-of-the-art results on all datasets. We compare against both 3D-only (top) and 2D–3D fusion methods (bottom). : Uses a smaller point patch size and relative positional encoding.
@@
Table 2: Outdoor semantic segmentation results (mIoU). We obtain state-of-the-art results on datasets with full camera coverage: nuScenes and Waymo. We compare against both 3D-only (top) and 2D–3D fusion methods (bottom).
Injection.  DITR achieves state-of-the-art semantic segmentation results across both indoor and outdoor datasets as shown in Tabs. 1 and 2. In particular, we outperform all previous 2D–3D fusion approaches, even though these methods either focus on indoor or outdoor settings. On the ScanNet200 hidden test set, we see an improvement of $\texttt{+}7.1$ mIoU, showing that image feature injection is especially beneficial for harder segmentation tasks with many classes. Among the outdoor datasets, only on the SemanticKITTI dataset DITR falls slightly behind the top-performing methods. Given that this dataset provides only a single front-facing camera, moderate improvements are to be expected. However, nuScenes stands out with a $\texttt{+}4.3$ mIoU improvement on the validation set. This could be caused by the fact that the nuScenes point clouds are sparser than those of the other datasets, thus resulting in a stronger boost from the dense image features.
@@
Table 3: Indoor distillation results (mIoU). We evaluate models pretrained and fine-tuned on the same datasets, as well as one pretrained on ScanNet and Structured3D. All models outperform the randomly initialized PTv3 baseline. : Uses a smaller point patch size and relative positional encoding.
@@
Distillation.  For distillation pretraining, we explore two settings: (1) pretraining on individual datasets and fine-tuning on the same dataset to show the effectiveness of VFM features as distillation targets, and (2) joint pretraining on multiple datasets. In the indoor multi-dataset distillation case, we jointly pretrain on ScanNet and Structured3D [67] and for the outdoor case we use nuScenes, SemanticKITTI, and Waymo. We exclude S3DIS from pretraining due to its smaller size and instead use it to assess generalization on an unseen dataset. The results are presented in Tab. 3 and Tab. 4 as D-DITR. Single-dataset distillation yields consistent improvements over the PTv3 baseline, demonstrating the effectiveness of DINOv2 distillation. The use of multi-dataset distillation further improves the results. In indoor datasets, distillation pretraining significantly improves the segmentation performance, with gains of $\texttt{+}2.4$ mIoU on ScanNet, $\texttt{+}2.3$ mIoU on ScanNet200 and $\texttt{+}2.9$ mIoU on S3DIS over the state-of-the-art 3D baseline. The improvement on S3DIS is particularly interesting, indicating that the distilled features can generalize to unseen datasets. For outdoor datasets, D-DITR also produces consistent improvements over the baseline. In particular, the distillation approach yields notable gains on SemanticKITTI, and even surpasses the injected DITR model (see Tab. 2). This indicates that even when images are not available, or compute constraints do not allow for the computation of DINOv2 features during inference, D-DITR can utilize these strong 2D features during pretraining for enhanced segmentation.
@@
Fig. 3 shows feature predictions of D-DITR prior to fine-tuning, on several indoor and outdoor scenes. When compared to the ground-truth segmentations, objects are clearly separable by color, even when visualizing only the first three principal components of the high-dimensional feature space. This further indicates the value of DINOv2 features for 3D semantic segmentation.
@@
Table 5: ScanNet limited reconstructions benchmark using the MinkUNet [7] 3D backbone. All methods perform pretraining on unannotated raw data, followed by segmentation fine-tuning on fixed labeled subsets of ScanNet. The baseline is a MinkUNet trained from scratch.
ScanNet Data Efficient Benchmark.  To demonstrate that D-DITR is also effective in label-scarce settings, we use the ScanNet limited reconstructions benchmark and show results for different percentages of available data during fine-tuning in Tab. 5. For this setting, we use MinkUNet [7] as the student model to compare with previous unsupervised approaches in a consistent setting. Similar to CSC [16], we use additional unlabeled images. D-DITR outperforms all previous unsupervised pretraining approaches in all settings. The gap is even larger when a small percentage of data is available during fine-tuning. Overall, this experiment shows that this distillation approach can be applied to other commonly used 3D backbones, providing a strong supervision signal compared to raw 3D data.
@@
Table 7: Number of images and image selection strategy during inference. We compare two inference-time image selection methods: random selection and temporally-equidistant frame selection (eq.dist.). For ScanNet200, we compare models trained on 6 (mIoU6) or 10 random images (mIoU10). For nuScenes, the model is trained using all six camera views. During inference, we use different subsets of cameras: no camera; the front camera alone; the front, back-left, and back-right cameras; all six cameras.
Visibility vs. Performance.  To better understand the source of the performance improvement, in Tab. 6, we compare the performance of DITR and the baseline on visible and invisible points separately. DITR shows a notable improvement on visible points for both the ScanNet200 and SemanticKITTI datasets. Moreover, it still outperforms the PTv3 baseline even on invisible points. The substantial improvement on ScanNet200 suggests that the model captures global context from the 2D features, aiding the segmentation of invisible regions.
@@
In Tab. 7, we analyze the impact of varying the number of images and the image selection strategy during inference while using the same trained model. For ScanNet200, we evaluate two models, one trained with 6 images per scene and another with 10 to evaluate the impact of the number of training images. For nuScenes, we only evaluate a model trained on all available images. On ScanNet200, the model trained with 6 images benefits from seeing more images during inference than during training and performs better than the one trained with 10 images when fewer images are used during inference. Furthermore, both models demonstrate robustness to seeing fewer images during inference, with performance degrading gracefully as they are trained with randomly selected images. However, on nuScenes, the model always receives the full surrounding visual context during training. As a result, the performance degrades significantly when frames are missing during inference. Finally, on ScanNet200, we also show the effectiveness of using temporally-equidistant images during inference, resulting in higher point coverage and consistently outperforming the random image selection.
Resource Usage.  We assess the runtime impact of DITR, compared to the PTv3 baseline without additional fusion of 2D features. On ScanNet200, using a ViT-L DINOv2 backbone and 10 camera views, training time increases from 12 to 15 hours using two H100 GPUs. The average per-scene inference latency increases from $41$ to $76\text{\,}\mathrm{ms}$, while the required GPU memory increases from $1.4$ to $5.6\text{\,}\mathrm{GiB}$ (using a single H100 GPU for inference). Although this is a limitation of DITR, we note that these inference penalties do not apply to D-DITR as the architecture remains unchanged.
4.4 Ablation Study
Injection point ScanNet200 nuScenes
None  (i.e. PTv3 [58]) 35.2 79.9
Before segmentation head 37.6 82.5
Before first encoder layer 40.1 82.3
Decoder (last block) 40.1 82.7
Decoder (all blocks)  (i.e. DITR) 41.2 83.1
Class tokens only 37.4 79.4
Table 8: Comparison of different injection points. We add the image features into the PTv3 model at different points throughout the network. All experiments use the DINOv2-L backbone.
Injection Location Ablation.  Tab. 8 compares different methods to inject image features into the 3D backbone. By default, we inject the DINOv2 features into all decoder blocks, yielding the strongest performance across the board. Injecting the features only at the last layer of the decoder is still a fairly strong option, albeit slightly worse. We then look at two other extremes: only injecting them directly prior to the segmentation head, or directly before the first encoder layer, after the points have been projected into a higher dimension. While both configurations improve the performance compared to the baseline, the late fusion approach yields considerably smaller gains on ScanNet200. As the primary computational overhead stems from the computation of VFM features, the location and frequency of injection into the backbone have negligible additional impact on efficiency. Thus, we recommend adding 2D features to all decoder layers.
Tab. 6 showed that DITR even achieves significant improvements on invisible points, suggesting that it might leverage additional features as a kind of global context. To assess this, we also experiment with injecting the class tokens for each image directly into the attention blocks of PTv3. While this does yield a small boost in indoor scenarios, it makes almost no difference in the outdoor case, and using the patch features yields considerably better results. Therefore, we did not further pursue class token injection.
Image backbone ScanNet200 nuScenes
Pretraining Model
None 35.2 80.4
IN21k [49] ViT-L 38.2 80.2
AIMv2 [12] ViT-L 39.1 82.8
SigLIP 2 [54] ViT-g 38.1 83.6
DINOv2 [33] ViT-S 38.2 82.8
DINOv2 [33] ViT-B 40.7 83.0
DINOv2 [33] ViT-L 41.2 83.1
DINOv2 [33] ViT-g 40.8 84.2
Image Backbone Ablation.  To assess the impact of the image backbone used to extract 2D features, we experiment with various commonly used image backbones with frozen weights as shown in Tab. 9. When comparing different DINOv2-pretrained ViT models, there is a consistent trend that larger models perform better. We also compare DINOv2 with other pretraining approaches. A standard ViT-L model pretrained on IN21k [9, 49] shows a modest performance improvement in indoor scenarios compared to the DINOv2 baseline, but reduces performance on the outdoor nuScenes dataset. We further evaluate two alternative foundation models, SigLIP 2 [54] and AIMv2 [12], both of which consistently improve performance in both indoor and outdoor settings. These results suggest that strong pretraining, as used in foundation models, is key to achieving consistent gains across diverse environments. They also indicate that our conclusions are not specific to DINOv2 and are likely to hold with future, even more powerful VFMs, which could serve as drop-in replacements within our framework.
5 Conclusion
Even though 3D foundation models might arise in the future, we argue that the DINO in the room is here to stay.Despite the success of 2D foundation models such as DINOv2, comparable models have yet to emerge in 3D, likely due to the scarcity of large-scale datasets. Therefore, we explore how existing 2D foundation models can be leveraged to enhance 3D segmentation performance. While 2D–3D fusion has been an active area of research, the potential of 2D VFMs remains relatively underexplored. With DITR, we demonstrate that even a straightforward injection of frozen VFM features into a 3D backbone can yield significant performance gains. Moreover, this approach is broadly compatible with both existing and most likely also future 3D backbones and VFMs. Even when images are unavailable during inference, our D-DITR distillation pretraining strategy can still provide substantial improvements. Notably, both DITR and D-DITR require only unlabeled images, making them readily applicable in most scenarios. Thus, we strongly advocate the use of 2D VFMs for 3D scene understanding whenever possible and hope that our findings will encourage more research in this direction.
Appendix A Implementation Details
Indoor Semantic Segmentation.
Following PTv3, the model is trained with a batch size of 12 for 800 epochs on ScanNet [8] and ScanNet200 [44], and for 3000 epochs on the smaller S3DIS [1] dataset. Optimization is performed using the AdamW [30] optimizer, with a 1cycle [48] learning rate schedule. The maximum learning rate is set to $6\text{\times}{10}^{-3}$ with the learning rate peaking at 5 % of the training process. Point cloud augmentations include random rotation, random dropout of points, random scaling, random flipping, elastic distortion, and color jittering. Grid sampling is applied with a grid size of $2\text{\,}\mathrm{cm}$.
We only inject features for points that are visible in the selected images. To prevent injecting features for occluded points, we compare the depth of each projected point with the raw depth value from the RGB-D sensor and exclude points where the depth difference exceeds a small margin of error. Additionally, to ensure that image features consistently are assigned to points at similar distances, points that are closer than $1\text{\,}\mathrm{m}$ or farther than $4\text{\,}\mathrm{m}$ are filtered out. Given that the 3D scene is captured from multiple camera angles with overlapping regions, most points fall in this range in at least one image.
Outdoor Semantic Segmentation.
Appendix B Qualitative Results

### v2

@@ 3.2 Distillation
While injecting DINOv2 features directly into the 3D backbone can significantly boost segmentation performance, relying on calibrated images at inference time may be restrictive in certain real-world scenarios. To address situations where only 3D data can be used at test time, we propose a distillation scheme (D-DITR) to transfer 2D knowledge into a pure 3D model as a pretraining step.
@@ 4.1 Datasets
We evaluate our method on well-established 3D indoor and outdoor segmentation benchmarks. For indoor segmentation, we select ScanNet [10], ScanNet200 [48], ScanNet++ v2 (SN++) [69], and S3DIS [1]. All four datasets consist of colored point clouds and corresponding RGB-D frames. For outdoor segmentation, we select nuScenes [6], SemanticKITTI (Sem.KITTI) [5], and Waymo [55].
@@ 4.2 Implementation Details
We use PTv3 [62] as our 3D backbone, as it is the state-of-the-art model for 3D semantic segmentation across both indoor and outdoor datasets. Unless stated otherwise, we retain the default architecture and hyperparameters of PTv3 to ensure fair comparison. Additional details are provided in the supplementary material. For all datasets, we adhere to standard evaluation protocols and metrics. Reproducing PTv3 results on the S3DIS [^1] https://github.com/Pointcept/Pointcept/issues/154 and SemanticKITTI [^2] https://github.com/Pointcept/Pointcept/issues/186 datasets has been notoriously hard for the community. Therefore, we also report our reproduced results as a reference.
@@
During both training and inference on outdoor scenes, we use all available camera views: one for SemanticKITTI, five for Waymo, and six for nuScenes. These camera views appear in a consistent orientation with respect to the LiDAR sensor and can be assumed always to be present. For the indoor datasets, the number of RGB-D frames per scene can vary tremendously and always surpasses the number of frames we can effectively process with the available GPU memory. Therefore, unless specified otherwise, we select 10 uniformly sampled, temporally equidistant frames from each RGB-D video during inference, while randomly sampling 10 views per scene during training to enhance data diversity. Since DINOv2 is pretrained at a maximum resolution of $518\times 518$, we resize input images to maintain a similar number of patches while preserving aspect ratio [59].
4.3 Main Results
Method ScanNet ScanNet200 S3DIS SN++
Val Test Val Test Area5 Test
@@ 4.3 Main Results
DITR 80.5 79.7 41.2 44.9 74.1 52.5
Table 1: Indoor semantic segmentation results (mIoU). DITR significantly outperforms PTv3 on all datasets and obtains state-of-the-art results on three of them. We compare against both 3D-only (top) and 2D–3D fusion methods (bottom). Uses a smaller point patch size and relative positional encoding. Latest results from the challenge in combination with multi-dataset fine-tuning. Requires 2D segmentation labels.
@@
Table 2: Outdoor semantic segmentation results (mIoU). DITR obtains state-of-the-art results on datasets with full camera coverage: nuScenes and Waymo. We compare against both 3D-only (top) and 2D–3D fusion methods (bottom).
Injection.  Tabs. 1 and 2 show that DITR significantly outperforms the reproduced PTv3 baseline [62] by injecting DINOv2 image features into the PTv3 backbone. These improvements are consistent across both indoor and outdoor benchmarks. We find that the performance improvement is particularly large on the hidden test set of ScanNet200, where we observe a gain of $\texttt{+}7.1$ mIoU. This shows that image feature injection is especially beneficial for more complex segmentation tasks with many categories. The only dataset where we observe only moderate gains is the Sem.KITTI dataset. However, this is to be expected because this dataset provides images from only a single front-facing camera, covering only a fraction of the 3D points, limiting the number of points for which 2D injection can be applied.
Compared to existing state-of-the-art methods, we note that DITR not only outperforms all previous 2D–3D fusion methods, but also the very recent Sonata [61] on most datasets. The improved performance compared to existing fusion methods is especially noteworthy because these methods only focus on either indoor or outdoor segmentation, while DITR is effective in both settings, and because some of them require 2D segmentation labels, which DITR does not. The fact that DITR surpasses Sonata, on the other hand, is striking because Sonata uses a PTv3 backbone that is three times larger than DITR’s, and because it uses an expanded 3D training corpus. Interestingly, we observe in Tab. 6 that DITR still outperforms Sonata even when using the small DINOv2-S backbone, with DITR’s model size—including a frozen DINOv2—amounting to only half of that of Sonata. This reinforces our main message: given the relatively limited availability of 3D data, 2D VFMs are crucial for advancing 3D segmentation, and even the inexpensive small variants can be highly effective.
@@
Table 3: Indoor distillation results (mIoU). We evaluate models pretrained and fine-tuned on the same datasets, as well as one pretrained on ScanNet and Structured3D jointly. All models outperform the randomly initialized PTv3 baseline. Uses a smaller point patch size and relative positional encoding.
@@
Distillation.  For distillation pretraining, we explore two settings: (1) pretraining on individual datasets and fine-tuning on the same dataset to show the effectiveness of VFM features as distillation targets, and (2) joint pretraining on multiple datasets. In the indoor multi-dataset distillation case, we jointly pretrain on ScanNet and Structured3D [74] and for the outdoor case we use nuScenes, SemanticKITTI, and Waymo. We exclude S3DIS from pretraining due to its smaller size and instead use it to assess generalization on an unseen dataset. The results are presented in Tab. 3 and Tab. 4 as D-DITR. We find that single-dataset distillation yields consistent improvements over the PTv3 baseline, demonstrating the effectiveness of DINOv2 distillation. The use of multi-dataset distillation further improves the results. In indoor datasets, distillation pretraining significantly improves the segmentation performance, with gains of $\texttt{+}2.4$ mIoU on ScanNet, $\texttt{+}2.3$ mIoU on ScanNet200 and $\texttt{+}2.9$ mIoU on S3DIS over the state-of-the-art 3D baseline. The improvement on S3DIS is particularly interesting, indicating that the distilled features can generalize to unseen datasets. For outdoor datasets, D-DITR also produces consistent improvements over the baseline. In particular, the distillation approach yields notable gains on SemanticKITTI, and even surpasses the injected DITR model (see Tab. 2). This indicates that even when images are not available, or compute constraints do not allow for the computation of DINOv2 features during inference, D-DITR can utilize the strong 2D features that it sees during pretraining for enhanced segmentation during inference.
@@
Fig. 3 visualizes feature predictions by D-DITR prior to fine-tuning. When compared to the ground-truth segmentation labels, objects are clearly separable by color, even when visualizing only the first three principal components of the high-dimensional feature space. This further indicates the value of 2D VFM features for 3D semantic segmentation.
4.4 Ablation Study
Injection method ScanNet200 nuScenes
None  (i.e., PTv3 [62]) 35.2 79.9
2D image feature injection
Before 3D encoder [23, 46] 40.1 82.3
Between two 3D encoder–decoders [67] 40.4 82.2
In 3D decoder (all blocks) (i.e., DITR) 41.2 83.1
In 3D decoder (last block only) 40.1 82.5
After 3D decoder 37.6 82.5
Other
[cls] tokens in 3D decoder (all blocks) 37.4 79.4
Table 5: Comparison of different 2D–3D injection methods. We compare different methods to inject 2D features, ordered from early to late injection. All experiments use DINOv2-L.
2D–3D Injection Method.  As described in Sec. 3.1, DITR injects 2D image features into all blocks of the 3D decoder. In Tab. 5, we compare this approach to several alternatives, including the 2D–3D injection methods that are employed by existing state-of-the-art 2D–3D fusion methods for indoor 3D segmentation that do not require 2D labels [23, 46, 67]. The results show that early injection—i.e., before the 3D encoder—as employed by MVPNet [23] and DVA [46], is better than no injection, but that it underperforms DITR’s intermediate injection in all decoder blocks. The same applies to DMF-Net’s approach [67] of having separate 3D encoder–decoders for feature extraction and 3D segmentation, and injecting features between these two encoder–decoders. Finally, we also evaluate late injection methods, which either inject 2D features only in the last decoder block or after the last decoder block, but we find that these also perform worse than DITR’s default injection. These results highlight the importance of informing the 3D model about 2D features at multiple stages of the network, as done by DITR’s 2D–3D injection approach.
In addition to the injection of 2D image features, we also experiment with injecting the [cls] tokens, generated by the 2D VFM, as additional tokens into the self-attention operations of the 3D decoder. This is inspired by the observation in Tab. 8 that DITR even achieves significant improvements on invisible points, suggesting that it might leverage additional features as a kind of global context, which could be contained by the [cls] token as well. However, we find that injecting the [cls] tokens only yields a small boost on indoor scenarios and does not impact outdoor performance. Therefore, we did not further pursue this direction.
Image backbone ScanNet200 nuScenes
Pretraining Model
35.2 80.4
IN21k [54] ViT-L 38.2 80.2
AIMv2 [14] ViT-L 39.1 82.8
SigLIP 2 [59] ViT-g 38.1 83.6
DINOv2 [36] ViT-S 38.2 82.8
DINOv2 [36] ViT-B 40.7 83.0
DINOv2 [36] ViT-L 41.2 83.1
DINOv2 [36] ViT-g 40.8 84.2
DINOv3 [52] ViT-L 42.3 83.9
Image Backbone Ablation.  To assess the impact of the image backbone used to extract 2D features, we experiment with various commonly used image backbones with frozen weights as shown in Tab. 6. When comparing different DINOv2-pretrained ViT models, there is a consistent trend that larger models perform better. Comparing DINOv2 with other pretraining approach, we find that a standard ViT-L model pretrained on IN21k [11, 54] shows a modest performance improvement in indoor scenarios compared to the no-injection baseline, but reduces performance on the outdoor nuScenes dataset. SigLIP 2 [59] and AIMv2 [14], on the other hand, consistently outperform the baseline in both indoor and outdoor settings, but perform worse than DINOv2. Finally, DINOv3 [52], which was released only days ago, performs even better than DINOv2 with ViT-L, and achieves the absolute best ScanNet200 performance. These results show that strong pretraining, as used in foundation models, is key to achieving consistent gains across diverse environments. They also suggest that our conclusions are not specific to DINOv2 but hold with very recent, more powerful VFMs like DINOv3 as well, and that these models are simply drop-in replacements in our framework.
4.5 Additional Analyses
@@
Table 7: ScanNet “limited reconstructions” benchmark with the MinkUNet [9] 3D backbone. All methods perform pretraining on unannotated raw data, followed by segmentation fine-tuning on fixed labeled subsets of ScanNet. For fair comparison, all methods use a MinkUNet backbone (not PTv3).
ScanNet Data Efficient Benchmark.  To demonstrate that D-DITR is also effective in label-scarce settings, we use the ScanNet “limited reconstructions” benchmark and show results for different percentages of available data during fine-tuning in Tab. 7. For this setting, we use MinkUNet [9] as the student model to compare with previous unsupervised approaches in a consistent setting. Similar to CSC [18], we use additional unlabeled images. D-DITR outperforms all previous unsupervised pretraining approaches in all settings. The gap is even larger when a small percentage of data is available during fine-tuning. Overall, this experiment shows that our distillation approach can be applied to other commonly used 3D backbones and that it provides a strong supervision signal compared to raw 3D data.
@@
Visibility vs. Performance.  To better understand the source of the performance improvement, in Tab. 8, we compare the performance of DITR and the baseline on visible and invisible points separately. DITR shows a notable improvement on visible points for both the ScanNet200 and SemanticKITTI datasets. Moreover, it still outperforms the PTv3 baseline even on invisible points. The large improvement on ScanNet200 suggests that the model captures global context from the 2D features, aiding the segmentation of invisible regions.
Resource Usage.  We assess the runtime impact of DITR, compared to the PTv3 baseline without additional injection of 2D features. On ScanNet200, using a ViT-L DINOv2 backbone and 10 camera views, training time increases from 12 to 15 hours using two H100 GPUs. The average per-scene inference latency increases from $41$ to $76\text{\,}\mathrm{ms}$, while the required GPU memory increases from $1.4$ to $5.6\text{\,}\mathrm{GiB}$ (using a single H100 GPU for inference). Although this is a limitation of DITR, we note that these inference penalties do not apply to D-DITR as the architecture remains unchanged in that setting.
5 Conclusion
With this work, to the best of our knowledge, we are the first to demonstrate that the rich semantic features of 2D VFMs, like DINOv2, can be leveraged to advance 3D segmentation performance by large margins. First, with DITR, we demonstrate that injecting frozen 2D VFM features into a 3D model’s decoder yields significant performance gains, achieving new state-of-the-art results. Second, we show that 2D VFMs can enable substantial improvements even when images are unavailable during inference, by using our D-DITR distillation strategy to pretrain a 3D backbone. Notably, both DITR and D-DITR require only unlabeled images and are not bound by the choice of VFM, making them readily and generally applicable. In conclusion, given our promising results, we strongly advocate the use of 2D VFMs for 3D scene understanding whenever possible.
Appendix A Implementation Details
Indoor Semantic Segmentation.
Following PTv3, we train the model with a batch size of 12 for 800 epochs on ScanNet [10] and ScanNet200 [48], and for 3000 epochs on the smaller S3DIS [1] dataset. For ScanNet++ [69], we again follow PTv3, dividing the large point clouds into overlapping $6\text{\,}\mathrm{m}$ by $6\text{\,}\mathrm{m}$ tiles with a stride of $3\text{\,}\mathrm{m}$, and training for 400 epochs. Optimization is performed using the AdamW [33] optimizer, with a 1cycle [53] learning rate schedule. The maximum learning rate is set to $6\text{\times}{10}^{-3}$ with the learning rate peaking at 5 % of the training process. Point cloud augmentations include random rotation, random dropout of points, random scaling, random flipping, elastic distortion, and color jittering. Grid sampling is applied with a grid size of $2\text{\,}\mathrm{cm}$.
We only inject features for points that are visible in the selected images. To prevent injecting features for occluded points, we compare the depth of each projected point with the raw depth value from the RGB-D sensor and exclude points where the depth difference exceeds a small margin of error ($20\text{\,}\mathrm{cm}$). Additionally, to ensure that image features are consistently assigned to points at similar distances, points that are closer than $1\text{\,}\mathrm{m}$ or farther than $4\text{\,}\mathrm{m}$ are filtered out. Given that the 3D scene is captured from multiple camera angles with overlapping regions, most points fall in this range in at least one image.
Outdoor Semantic Segmentation.
Appendix B Impact of Image Selection Strategy
@@
Table 9: Impact of image selection strategy during inference. We compare two inference-time image selection methods: random selection and temporally-equidistant frame selection (eq.dist.). For ScanNet200, we compare models trained on 6 (mIoU6) or 10 random images (mIoU10). For nuScenes, the model is trained using all six camera views. During inference, we use different subsets of cameras: no camera; the front camera alone; the front, rear-left, and rear-right cameras; all six cameras.
@@
In Tab. 9, we analyze the impact of varying the number of images and the image selection strategy during inference while using the same trained model. For ScanNet200, we evaluate two models, one trained with 6 images per scene (mIoU6) and another with 10 (mIoU10) to evaluate the impact of the number of training images. For nuScenes, we only evaluate a model trained on all available images. We find that, on ScanNet200, the model trained with 6 images benefits from seeing more images during inference than during training, and performs better than the one trained with 10 images when fewer images are used during inference. Furthermore, both models demonstrate robustness to seeing fewer images during inference, with performance degrading gracefully as they are trained with randomly selected images. However, on nuScenes, the model always receives the full surrounding visual context during training. As a result, the performance degrades significantly when frames are missing during inference. Finally, on ScanNet200, we also show the effectiveness of using temporally-equidistant images during inference, resulting in higher point coverage and consistently outperforming the alternative of selecting frames randomly.
Appendix C Qualitative Results